OpenStudy (anonymous):

can you help me to my problems, can you show me the proper way to solve this one to east to understant? y =(x + 1) (2x - 1) diff calculus. thanks

OpenStudy (anonymous):

find the derivative again? multiply out first

OpenStudy (anonymous):

yes sir, :((

OpenStudy (anonymous):

its hard to me to find the derivative :((

OpenStudy (anonymous):

yeah i know lets multiply first then it will be not too hard at all what you get when you multiply \[(x+1)(2x-1)\]

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

pardon?

OpenStudy (anonymous):

you know what i mean by "multiply"? remove the parentheses using the distributive law or what criminal math teacher call "foil"

OpenStudy (anonymous):

so if we multiply?

OpenStudy (anonymous):

yes multiply what do you get ?

OpenStudy (anonymous):

2? (x +1 ) (2x -1)

OpenStudy (anonymous):

correct me if im wrong

OpenStudy (anonymous):

i guess it is not clear what i am saying

OpenStudy (anonymous):

sorry :(

OpenStudy (anonymous):

for example \[(x+5)(x+4)=x^2+5x+4x+20=x^2+9x+20\]

OpenStudy (anonymous):

ahh so it will happen is 1x + 2x^2?

OpenStudy (anonymous):

er no

OpenStudy (here_to_help15):

Sorry to intrude but to be correct she said easy to understand lol and if i was new to this, i would think what you wrote was Complex Calculus problems

OpenStudy (here_to_help15):

I am just saying @satellite73

OpenStudy (anonymous):

@Here_to_Help15 if you were going the find the derivative of \[y=(x +1 ) (2x -1)\] i am sure step one would be to multiply right? not for example use the product rule

OpenStudy (here_to_help15):

Yes i believe so

OpenStudy (anonymous):

without this skill (not to be rude) it is silly to be learning calculus

OpenStudy (here_to_help15):

lol i agree :D

OpenStudy (anonymous):

lets begin again. sir satellite

OpenStudy (here_to_help15):

See here if you are going to help her your going to lower your high intense brain down a notch and explain step by step

OpenStudy (anonymous):

1) multiply out once you have done so, we can use the 2) power rule \(\frac{d}{dx}[x^n]=nx^{n-1}\) and it will be brief

OpenStudy (anonymous):

i am not doing anything high powered at all, just trying to say "multiply" i have no other way of saying it distribute foil whatever

OpenStudy (here_to_help15):

Oh quick question ik this happens to be irrelevant but why does ppls reply say its from yesterday like yours for instance says yesterday?

OpenStudy (here_to_help15):

Is it because OS is lagging or my internet connection is bad?

OpenStudy (anonymous):

sir @satellite73 if we use the power rule what wil gonna happen to our solving?

OpenStudy (anonymous):

can you show me the proper way to solve again sir to easy to understand?

OpenStudy (anonymous):

\[y=(x +1 ) (2x -1)\] \[y=2x^2+2x-x-1\\ y=2x^2+x-1\] is a start

OpenStudy (anonymous):

so the y+ 2x^2 + 2 -1? the final?

OpenStudy (anonymous):

no that is what you get before you take the derivative using the power rule you get \[y'=4x+1\]

OpenStudy (anonymous):

so sir that the final right?

OpenStudy (anonymous):

thank you sir @satellite73 ill appreciate it for helping me :))